\begin{thebibliography}{abc99xyz}
\bibitem{Achdou 1} Y.Achdou, M-K. Dao, O. Ley, and N. Tchou. A class of infinite horizon mean field games on networks. Netw. Heterog. Media, 14(3):537–566, 2019.
\bibitem{Achdou 2}  Y.Achdou, M-K. Dao, O. Ley, and N. Tchou.. Finite horizon mean field games on networks. Calc. Var. Partial Differential Equations, 59(5):Paper No. 157, 34, 2020.
\bibitem{Barles Neuman 1} G. Barles. Fully nonlinear Neumann type boundary conditions for second-order elliptic and parabolic equations. J. Differ. Equ. 106, 90–106, 1993.
\bibitem{Barles Neuman 2} G. Barles. Nonlinear Neumann boundary conditions for quasilinear degenerate elliptic equations and applications. J. Differ. Equ. 154, 91–224, 1999.
\bibitem{control 1}G. Barles, A. Briani and E. Chasseigne. A Bellman approach for two-domains optimal control
problems in $\R^N$. ESAIM Control Optim. Calc. Var. 19, no. 3, 2013.
\bibitem{control 2} G. Barles, A. Briani and E. Chasseigne. A Bellman approach for regional optimal control problems
in $\R^N$. SIAM J. Control Optim. 52 , no. 3, 2014.
\bibitem{control 3} G. Barles and E. Chasseigne. (Almost) everything you always wanted to know about deterministic
control problems in stratified domains. Netw. Heterog. Media 10, no. 4, 2015.
\bibitem{Barles book} G. Barles and E. Chasseigne. On Modern Approaches of Hamilton-Jacobi Equations and Control Problems with Discontinuities. A Guide to Theory, Applications, and Some Open Problems. Progress in nonlinear Differential Equations and Their Applications, (104), 2024.
\bibitem{Barles semi linear} G. Barles, O. Ley and E. Topp. Degenerate Elliptic PDEs on a Network with Kirchhoff Conditions. ArXiv:2509.12848.
\bibitem{control 4} A. Bressan and Y. Hong. Optimal control problems on stratified domains. Netw. Heterog. Media.,
2, 2007.
\bibitem{Camilli 1} F. Camilli and C. Marchi. Stationary mean field games systems defined on networks. SIAM J. Control Optim., 54(2):1085–1103, 2016.
\bibitem{Camilli 2} F. Camilli, C. Marchi, and D. Schieborn. The vanishing viscosity limit for Hamilton-Jacobi equations on networks. J. Differential Equations, 254(10):4122-4143, 2013.
\bibitem{Carda junction -1} P. Cardaliaguet. A note on contractive semi-groups on a 1:1 junction for scalar conservation laws and Hamilton-Jacobi equations. ArXiv:2411.12326.
\bibitem{Carda junction 0} P. Cardaliaguet. A microscopic derivation of a traffic flow model on a junction with two entry lines. ESAIM: M2AN 59 (2025) 2021–2054, 2025.
\bibitem{Carda junction 1} P. Cardaliaguet, N. Forcadel, T. Girard and R.Monneau. Conservation laws and Hamilton-Jacobi equations on a junction: The convex case. Discrete and Continuous Dynamical Systems
Vol. 44, No. 12, pp. 3920-3961, 2024.
\bibitem{Carda junction 2} P. Cardaliaguet and P. E. Souganidis. An optimal control problem of traffic flow on a junction. ESAIM: COCV
Volume 30, 88, 2024.
%\bibitem{User guide} M.G. Crandall, H. Ishii and P.L. Lions. User’s guide to viscosity solutions of second
order Partial Differential Equations. Bull. Amer. Soc., 27, 1992.
\bibitem{Daniel game} J.P. Daniel. A game interpretation of the Neumann problem for fully nonlinear parabolic and elliptic equations. ESAIM: Control, Optimisation and Calculus of Variations, (19)  1109–1165, 2013. 
\bibitem{control 5} C. Imbert and R. Monneau. Flux-limited solutions for quasi-convex Hamilton Jacobi equations on
networks preprint. Annales scientifiques de l'école 
normale supérieure, 50, fascicule 2, 2017.
\bibitem{control 6} C. Imbert and V. D. Nguyen. Generalized junction conditions for degenerate parabolic equations. Calculus of Variations and PDE’s, 56, 2017.
%\bibitem{EDO} A.Granas, R.Guenther and J.Lee. nonlinear boundary value problems for ordinary differential equations. Instytut Matematyczny Polskiej Akademi Nauk (Warszawa), 1985.
\bibitem{Hishi neumann} H. Ishii. Fully nonlinear oblique derivative problems for nonlinear second-order elliptic PDEs. Duke Math. J. 62, 633–661, 1991.
\bibitem{Lions neumann} P.L. Lions. Neumann type boundary conditions for Hamilton-Jacobi equations. Duke Math. J. 52 793–820 1985.
\bibitem{Lions Souganidis 1} P.L.Lions and P.Souganidis. Viscosity solutions for junctions: well-posedness and stability. Rend. Lincei Mat. Appl. 27, 2016. 
\bibitem{Lions Souganidis 2} P.L.Lions and P.Souganidis. Well-posedness for multi-dimensional junction problems with Kirchhoff-type conditions. Rend. Lincei Mat. Appl. 28, 2017.
%\bibitem{Lions Souganidis 3} P.L.Lions and P.Souganidis. Effective transmission conditions for second-order elliptic equations on
networks in the limit of thin domains. Compte rendus Mathematiques. Vol. 358, Issue 7, 2020.
\bibitem{Spider} M.Martinez and I.Ohavi. Martingale problem for a Walsh's spider diffusion with spinning measure selected from its own local time. Electronic Journal of Probability, Volume 30, paper no 22, 2025.
\bibitem{Spider 2} On spider diffusion with spinning measure selected from its own local time. Submitted 2025, ArXiv:2502.02754. In progress in Stochastic Processes and their Applications.
\bibitem{linear PDE} M.Martinez and I.Ohavi. Well-posedness of linear Parabolic partial differential equation posed on a
star-shaped network with local-time Kirchhoff’s boundary condition at the vertex. Journal of Mathematical Analysis and Applications, 537 (2), 128-294, 2024.
\bibitem{Ohavi visco} I.Ohavi. Comparison principle for Walsh's spider HJB equations with nonlinear local time Kirchhoff's
boundary transmission. Journal of Mathematical Analysis and Applications, 547 (2), 2025.
\bibitem{Ohavi PDE} I.Ohavi. Quasi-linear parabolic PDE posed on a network with nonlinear Neumann boundary condition at vertices. Journal of Mathematical Analysis and Applications. Volume 500, Issue 1, 2021.
\bibitem{Ohavi control} I.Ohavi. Stochastic scattering control of spider diffusion governed by an optimal probability diffraction measure selected from its own local time. Submitted 2025, ArXiv:2501.18057. In progress in SIAM Journal on Control and Optimization.
\bibitem{Sato} M.-H. Sato. Interface evolution with Neumann boundary condition. Adv. Math. Sci. Appl. 4, 249–264, 1994.
\bibitem{Below 3} J.Von Below. A maximum principle for semi linear parabolic network equations. In Differential equations with applications in biology, physics, and engineering (Leibnitz, 1989), volume 133 of Lecture Notes in Pure and Appl. Math., Dekker, New York, 1991. 
\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{amssymb,amsfonts,amsthm}
\usepackage{ulem}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{frcursive}
\usepackage[english,american]{babel}
\usepackage{enumerate}
\usepackage{graphicx}
\usepackage{tikz} 
\usepackage{setspace}
\usepackage{hyperref}
\usepackage{color}

\definecolor{Red}{cmyk}{0,1,1,0.2}

\usepackage{amsfonts,amsmath,layout}
\newcommand{\R}{\mathbb R}

\def\R{\mathbb R}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}

\def\1{{\bf 1}}

\def\ds{\displaystyle}
\newtheorem{Theorem}{Theorem}[section]
\newtheorem{Definition}[Theorem]{Definition}
\newtheorem{Proposition}[Theorem]{Proposition}

\newtheorem{Remark}[Theorem]{Remark}

\begin{document}
\title[Nonlinear PDE posed on networks]{Viscosity solutions posed on star-shaped network with Kirchhoff's boundary condition: well posedness}
\space
\space
\author[Isaac Ohavi]{Isaac Ohavi $^\aleph$
\\
$^\aleph$Hebrew University of Jerusalem, Department of  Mathematics and Statistics, Israël}
\address{ Hebrew University of Jerusalem, Department of Mathematics and Statistics, Israël}
\email{isaac.ohavi@mail.huji.ac.il \& isaac.ohavi@gmail.com}
\dedicatory{Version: \today}
\maketitle
\begin{abstract} The aim of this work is to establish the well-posedness of fully nonlinear partial differential equations (PDE) posed on a star-shaped network, having nonlinear Kirchhoff's boundary condition at the vertex, and possibly degenerate. We obtain a comparison theorem, for discontinuous viscosity solutions, following the recent ideas obtained in \cite{Ohavi visco} for second order problems, building test functions at the vertex solutions of Eikonal equations with well-designed coefficients. Another strong result obtained in this contribution is to show that any generalized Kirchhoff's viscosity solution introduced by Lions-Souganidis in \cite{Lions Souganidis 1}-\cite{Lions Souganidis 2} is indeed a Kirchhoff's viscosity solution. In other terms, the values of the Hamiltonians are not required at the vertex in the analysis of these types of PDE systems.
\end{abstract}
{\small \textbf{Key words:}  Discontinuous Hamilton-Jacobi-Bellman equations, Kirchhoff's boundary condition, Comparison principle, Viscosity solutions.}\\
{\small \textbf{AMS Class:} 35F21, 49L25, 35B51, 49L20.}
\section{Introduction}

We are given an integer number $I$ (with $I\ge 2$) and a star-shaped compact network: $$\displaystyle \mathcal{N}_R=\bigcup_{i=1}^I\mathcal{R}_i,$$ that consists of $I$ compact rays $\mathcal{R}_i\cong [0,R]$ ($R>0$) emanating from a junction point $\bf 0$. We set in the rest of this work $[I]=\{1,\ldots,I\}$.
More precisely, the system involved here is the following:
\label{sec intro}
\begin{eqnarray}\label{eq PDE 0}
&\begin{cases}
\textbf{Second order HJB equations on each ray :}\\ 
H_i\big(x,u_i(x),\partial_xu_i(x),\partial_x^2u_i(x)\big)=0,~~x\in(0,R),\\
\textbf{Nonlinear Kirchhoff's boundary condition at }\bf 0:\\
F\big(u(0),\partial_xu_1(0),\dots,\partial_xu_I(0)\big)=0,\\
\textbf{Dirichlet's boundary condition at }x=R:\\
u_i(R)=a_i,~~\forall i\in [I].
\end{cases}
\end{eqnarray}
Note that in all of this work, in order to simplify our study, we have assumed in our framework that all the rays $\mathcal{R}_i=[0,R]\times \{i\},~i\in[I]$, have the same length $R>0$. A more general setting could be treated with similar tools: one could for instance, consider more general networks and/or a mix of Kirchhoff's and Dirichlet boundary conditions at the vertexes.

In this contribution, our concern is to obtain the well-posedness of the system \eqref{eq PDE 0} in the second order possibly degenerate framework. We are able to deal with equations whose nature at the junctions ranges from first-order Hamilton Jacobi Equations to uniformly elliptic second-order. Our viscosity solutions will be assumed to be discontinuous, that is l.s.c. for super solutions (resp. u.s.c for sub solutions). We establish three main results, which are:\\
-any generalized Kirchhoff's viscosity solution introduced by Lions-Souganidis in \cite{Lions Souganidis 1}-\cite{Lions Souganidis 2}, is indeed a Kirchhoff's viscosity solution,\\
-a comparison theorem,\\
-existence and uniqueness of continuous viscosity solution for the system \eqref{eq PDE 0}.
%Because this contribution is already quite long, and presents new ideas within the framework of viscosity theory for networks problem, we have decided to push back in a
%subsequent work the second order case. Even if we believe that strictly elliptic case will behave like the HJB systems studied in \cite{Ohavi visco}, our concern in the next contribution is to adapt the ideas of this work, for the second order possible degenerate  case. Therein, we will also try to understand how the deterministic 'local-time' variable introduced in \cite{Ohavi visco}, can impact the degenerate case, and address also some results on vanishing limit viscosity solutions.

We follow the recent ideas obtained in \cite{Ohavi visco} for second order strictly elliptic problems, for the construction of test functions at the vertex. These test functions for super solutions (resp. sub solutions) denoted $\overline{\phi}$ (resp. $\underline{\phi}$), are indeed solutions of Eikonal equations system posed on a star-shaped network, with coefficients that may be viewed as a kind of envelope of all possible errors of the Hamiltonians, and well-chosen boundary conditions. The expressions of the test functions $(\overline{\phi},\underline{\phi})$ will be determinant to answer all the problems stated in this contribution. Let us explain.

First, these last test functions are constructed so that the signs of their Hamiltonians at the junction point imply that only the Kirchhoff's condition is verified at the vertex, in the definition of generalized Kirchhoff's solutions of Lions-Souganidis in \cite{Lions Souganidis 1}-\cite{Lions Souganidis 2}. Moreover, we will show that they have an optimal gradient, in the sense that: any test functions for a super/sub solution at $\bf 0$, have a gradient lower/upper bounded by the gradient of $(\overline{\phi}/\underline{\phi})$ at $\bf 0$. This then will imply that generalized Kirchhoff's viscosity solutions are indeed nothing more than Kirchhoff's viscosity solutions. This result appears as a major advance in the formulation of viscosity solutions posed on network. Beyond showing that a weak boundary condition is a strong one, it also allows to avoid all the discontinuities induced by the Hamiltonians at $\bf 0$. The discontinuities induced by the Hamiltonians at $\bf 0$ appeared as the main technical difficulty in the analysis of PDE systems of type \eqref{eq PDE 0} the last decade, since the doubling method variable fails in this case.
Finally, the last main result of this contribution will be to establish a comparison theorem using once again the expressions of $(\overline{\phi}/\underline{\phi})$. Note that existence will be performed using classical arguments of Perron's method.

We will first investigate the first order case in Section \ref{sec 1 ordre} for system \eqref{eq PDE 0}, to simplify the calculations and give the readers the best intuitions for the sake of clarity. We will see that, up to some modifications, our results can be extended easily to the second order possibly degenerate framework.
Notably, the results of this work do not require any convexity and interpretation with control theory. We use pure PDE arguments, and classical tools in the field of viscosity theory can generalize the main results of this contribution to the parabolic framework for system \eqref{eq PDE 0}.

Let us now give an account of the main works that have been done in literature for similar systems close to \eqref{eq PDE 0}.

We first refer to the main results contained in \cite{Ohavi visco} that have given us the seeds of the techniques used in this paper. Therein, the first result in literature that gives a comparison theorem for fully nonlinear HJB systems (strictly elliptic) of type \eqref{eq PDE 0} was proved, with a more general boundary condition called: \textit{nonlinear local time Kirchhoff's boundary condition}. Notably, the author uses an original technique consisting of introducing a local time variable that drives the system only at the junction point $\bf 0$. The main idea of the proof of the comparison theorem is to build test function with local time derivative that absorbs at $\bf 0$ the error term induced by the \textit{Kirchhoff's speed of the Hamiltonians}.
Also note that the  \textit{nonlinear local time Kirchhoff's boundary condition} has been used in  \cite{Ohavi visco} to show a comparison theorem for non degenerated systems of type \eqref{eq PDE 0}.
Thereafter, as a natural main direction, the results contained in \cite{Ohavi visco} allowed us to formulate a problem of stochastic scattering control for spider motion having a spinning measure selected from its own local time at $\bf 0$, introduced in \cite{Spider}-\cite{Spider 2}. Indeed, the authors show in \cite{Ohavi control} that the optimal diffraction of a spider at the vertex $\bf 0$, can be characterized with the aid of a value function solution of HJB system of type \eqref{eq PDE 0}, involving an optimal diffraction probability measure in the Kirchhoff's condition, selected from its own local time. Remark that the Kirchhoff's condition in \cite{Ohavi visco} is formulated in the strong sense, implying that the values of the Hamiltonians at $\bf 0$ are not required. We refer also to \cite{linear PDE} where linear systems having the \textit{nonlinear local time Kirchhoff's boundary condition} have been studied, and the results contained there.
For a recent monograph that presents some of the most recent developments in the study of Hamilton-Jacobi Equations and control problems with discontinuities, (excluding \cite{Ohavi visco} and \cite{linear PDE}), we refer to \cite{Barles book} (see Part III for the case of problems on networks). For quasi linear and semi linear systems of type \eqref{eq PDE 0} for which the solution is differentiable at $\bf 0$, we refer to \cite{Ohavi PDE} \cite{Camilli 1}, \cite{Camilli 2}, \cite{Achdou 1}, \cite{Achdou 2}, \cite{Barles semi linear}, \cite{Below 3}... 
In \cite{Lions Souganidis 1}, the authors introduce a notion of state-constraint viscosity solutions for one dimensional “junction” - type problems for first order Hamilton-Jacobi equations with non-convex coercive Hamiltonians and study its well-posedness and stability properties. On the other hand, in \cite{Lions Souganidis 2} the authors have studied multi-dimensional junction problems for first and second-order PDE with Kirchhoff-type
Neumann boundary conditions, showing that their generalized viscosity solutions are unique, but still with a vanishing viscosity at the vertex for the second order terms. In both of the last works, generalized Kirchhoff's boundary conditions are considered, and in the same time it is assumed that the viscosity solutions are Lipschitz continuous at $\bf 0$, that allows to avoid the doubling method variable and simplify the comparison theorem. Note that in \cite{Lions Souganidis 2}, it is proved that Flux-limiter solutions are generalized Kirchhoff's solutions in the sense of Imbert-Monneau in \cite{control 5}-\cite{control 6}, for an appropriate choice of the constant part appearing in the Kirchhoff's boundary condition.
For some references on literature based on the control (deterministic) theoretical interpretation, among them we can cite the long list of references on this topic with convex Hamiltonians, with: \cite{control 1}, \cite{control 2}, \cite{control 3}, \cite{control 4}, \cite{control 5}, \cite{control 6}. For recent works on systems of conservative laws posed on junctions, we refer also to \cite{Carda junction -1}, \cite{Carda junction 0}, \cite{Carda junction 1} and \cite{Carda junction 2} and the references therein. 
Finally, note that there are already a lot of comparison and existence results for viscosity solutions of second order PDEs with general Neumann type boundary conditions for classical problems. We refer for this to \cite{Barles Neuman 1}, \cite{Barles Neuman 2}, \cite{Daniel game},   \cite{Hishi neumann}, \cite{Lions neumann}, \cite{Sato} and references therein.

The paper is organized as follows: In Section \ref{sec : Notations}, we introduce the main notations-assumptions, and we state the main Theorems of this work: Theorem \ref{th : weak are strong}, Theorem \ref{th : compa syst} and Theorem \ref{th : existence}. To simplify the presentation of the proofs, the last three main results are proved in Section \ref{sec 1 ordre} in the case of first order problems involving systems of type \eqref{eq PDE 0}. Finally, we prove our main results for the second order case, possibly degenerate in Section \ref{sec 2 order}.
\section{Notations, Definitions and Main results} \label{sec : Notations}
We are given an integer number $I$ (with $I\ge 2$) and a star-shaped compact network: $$\displaystyle \mathcal{N}_R=\bigcup_{i=1}^I\mathcal{R}_i,~~[I]:=\{1,\ldots,I\},$$ that consists of $I$ compact rays $\mathcal{R}_i\cong [0,R]$ ($R>0$) emanating from a junction point $\bf 0$. The intersection of the $(\mathcal{R}_{i})_{i \in [I]}$ is called the junction point and is denoted by $\bf 0$.
We identify  all the points of $\mathcal{N}_R$ by the couples $(x,i)$ (with $i \in[I], x\in[0,R]$), such that we have: $(x,i)\in \mathcal{N}_R$ if and only if $x\in \mathcal{R}_{i}$, whereas the junction point ${\bf 0}$ is identified with the equivalent class $\{(0,i)~:~i\in [I]\}$.
We introduce the following data:
$$\textbf{Data:~} (\mathcal{D})~~\begin{cases}
\Big(Q_i \in \mathcal{C}\big([0,R]\times [0,+\infty)^3,\R\big)\Big)_{i\in[I]}\\
F\in \mathcal{C}\big(\R\times \R^I,\R\big)\\
\big(a_i\big)_{i\in\{1 \ldots I\}}\in \R^I,
\end{cases}.
$$
We assume that the data $(\mathcal{D})$ satisfy the following assumption:
$$\textbf{Assumption } (\mathcal{H})$$
a) The growth assumption for the nonlinear terms $(Q_i)_{i\in[I]}$ with respect to their second variable: there exists a strictly positive constant $\lambda>0$, independent of $i\in[I]$,
such that:
\begin{eqnarray*}
& \forall (x,u,v,p,X)\in [0,R]\times \R^4,~~\text{if}: v\leq u,~\text{then}~~\forall i\in[I]:\\
&Q_i(x,u,p,X)-Q_i(x,v,p,X)\ge \lambda(u-v).
\end{eqnarray*}
b) The classical assumption for the nonlinear terms $(Q_i)_{i\in[I]}$, compatible with the Ishii's matrix lemma, that reads:
\begin{eqnarray*}
 &\exists \omega \in  \mathcal{C}(\R_+,\R_+),~~\omega(0)=0,~~
 \forall \alpha>0,~~\forall (x,y,u,p,X,Y)\in [0,R]^2\times \R^4,~\text{such that:}\\
 &-3\alpha \begin{pmatrix}1&0\\
0&1
 \end{pmatrix}\leq\begin{pmatrix}X&0\\
 0&-Y
 \end{pmatrix}\leq 3\alpha \begin{pmatrix}1&-1\\
-1&1
 \end{pmatrix},~~\text{then}:\\
&\forall i\in[I]:~~Q_{i}(y,u,\alpha(x-y).Y)-Q_{i}(x,u,\alpha(x-y),X)\leq \omega\big(\alpha|x-y|^2+|x-y|\big).
\end{eqnarray*}
c) The nonlinear terms $(Q_i)_{i\in[I]}$ are elliptic possibly degenerate, that is: 
\begin{eqnarray*}
&\forall (x,u,p,X,Y)\in[0,R]\times\R^4,~~\text{if:}~~X\ge Y,~\text{then}~~\forall i\in[I]:\\
&Q_i(x,u,p,Y)\ge Q_i(x,u,p,X).  
\end{eqnarray*}
d) A Lipschitz growth condition with respect to the gradient of the nonlinear terms $(Q_i)_{i\in[I]}$:
\begin{eqnarray*}
&\forall (x,u,p,X)\in[0,R]\times\R^3,~~\text{if:}~~|u|\leq M,~~\text{and}~~|X|\leq K,~\text{then}~~\exists C_{M,K}>0,\\
&~~\text{s.t}~~\forall i\in[I]:~~|Q_i(x,u,p,X)|\leq C_{M,K}(1+|p|).  
\end{eqnarray*}
e) The nonlinear Kirchhoff's term $F$ is non increasing with respect to its first variable:
\begin{align*}
&\forall (u,v,p_1,\ldots p_I)\in \R^2 \times \R^I,~\text{if}~u\ge v,~~\text{then}:~~F(u,p_1,\ldots p_I)\leq F(v,p_1,\ldots p_I).
\end{align*}
f) The nonlinear Kirchhoff's term $F$ is strictly increasing with respect to its gradient, that reads:
\begin{align*}
&\forall (u,p_1,\ldots p_I,q_1,\ldots,q_I)\in \R\times \R^I\times \R^I,~~\exists \alpha >0,~~\text{s.t.}~~\text{if:}~~\forall i\in[I]~~p_i\ge q_i,~\text{then:}\\
&~~~~~~~~~~~~
F(u,p_1,\ldots p_I)-F(u,q_1,\ldots q_I)\ge \alpha \big(\sum_{i=1}^I p_i-q_i\big).
\end{align*}
Next, we introduce the set of lower semi-continuous $\mathcal{L}_{sc}(\mathcal{N}_R,\R)$ and upper semi-continuous $\mathcal{U}_{sc}(\mathcal{N}_R,\R)$ functions defined on the star-shaped network $\mathcal{N}_R$, and valued on $\R$: 
\begin{align*}
&\mathcal{L}_{sc}(\mathcal{N}_R,\R):=\Big\{~~f:\mathcal{N}_R\to \R,~~(x,i)\mapsto f_i(x)~~\text{ s.t}~~\forall i\in [I]~~x\mapsto f_i(x)~~\text{is lower}\\
&\hspace{3cm}\text{semi-continuous in}~ [0,R],~~\text{and}~~f({\bf 0})=\underset{j\in [I]}{\text{min}}f_j(0)~~\Big\},\\
&\mathcal{U}_{sc}(\mathcal{N}_R,\R):=\Big\{~~f:\mathcal{N}_R\to \R,~~(x,i)\mapsto f_i(x)~~\text{ s.t}~~\forall i\in [I]~~x\mapsto f_i(x)~~\text{is upper} \\
&\hspace{3cm}\text{semi-continuous in}~ [0,R],~~\text{and}~~f({\bf 0})=\underset{j\in [I]}{\text{max}}f_j(0)~~\Big\}.
\end{align*}
The last definitions of the sets $\mathcal{U}_{sc}(\mathcal{N}_R,\R)$ and $\mathcal{U}_{sc}(\mathcal{N}_R,\R)$, are the ones that remain compatible to define the semi-continuity on $\mathcal{N}_R$, if one considers the geodesic metric $d^{\mathcal{N}_R}$ defined on the star-shaped network $\mathcal{N}_R$ defined by:
\begin{equation*}
\forall \Big((x,i),(y,j)\Big)\in \mathcal{N}_R^2,~~d^\mathcal{N_R}\Big((x,i),(y,j) \Big)  := \left\{
\begin{array}{ccc}
 |x-y| & \mbox{if }  & i=j\;,\\ 
x+y & \mbox{if }  & {i\neq j},\;
\end{array}\right.
\end{equation*}
We denote also $\mathcal{L}_{sc}^b(\mathcal{N}_R,\R)$ (resp. $\mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$) the subsets of $\mathcal{L}_{sc}(\mathcal{N}_R,\R)$ (resp. $\mathcal{U}_{sc}(\mathcal{N}_R,\R)$) of bounded semi-continuous maps defined on $\mathcal{N}_R$.

Next, we give the definition of generalized viscous solution of system \eqref{eq PDE 0}, in the sense of Lions-Souganidis in \cite{Lions Souganidis 1}-\cite{Lions Souganidis 2}, called generalized viscous Kirchhoff's solution. We also define the strong version of the last solutions, in the spirit of Ohavi in \cite{Ohavi visco}, that we call viscous Kirchhoff's solution.
\begin{Definition} \label{def : weak sur/sous solutions}\text{Generalized viscous Kirchhoff's solutions: Lions-Souganidis \cite{Lions Souganidis 1}-\cite{Lions Souganidis 2}}:\\
a) We say that $u\in \mathcal{L}_{sc}^b(\mathcal{N}_R,\R)$ is generalized viscosity Kirchhoff's super solution of
\eqref{eq PDE 0}, if for all test function $\overline{\phi} \in \mathcal{C}^{2}\big(\mathcal{N}_R,\R\big)$, and for all local minimum point $(x_0,i_0)\in \mathcal{N}_R$ of $u-\overline{\phi}$, with $u_{i_0}(x_0)-\overline{\phi}_{i_0}(x_0)=0$, we have:
\begin{eqnarray*}
\begin{cases}
Q_{i_0}\big(x_0,u_{i_0}(x_0),\partial_x\overline{\phi}_{i_0}(x_0),\partial_x^2\overline{\phi}_{i_0}(x_0)\big)\ge 0,~\text{if}~x_0 \in (0,R),\\
\max\Big\{\underset{i\in[I]}{\max}\{Q_{i}\big(0,u(0),\partial_x\overline{\phi}_{i}(0),\partial_x^2\overline{\phi}_{i}(0)\big)\},-F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial \overline{\phi}_I(0)\big)\Big\}\ge 0,\\
\text{if}~x_0 = \bf 0.
\end{cases}
\end{eqnarray*}
b) We say that $u\in \mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$ is a generalized viscosity Kirchhoff's sub solution of
\eqref{eq PDE 0}, if for all test function $\overline{\phi} \in \mathcal{C}^{2}\big(\mathcal{N}_R,\R\big)$, and for all local maximum point $(x_0,i_0)\in \mathcal{N}_R$ of $u-\overline{\phi}$, with $u_{i_0}(x_0)-\overline{\phi}_{i_0}(x_0)=0$, we have:
\begin{eqnarray*}
\begin{cases}
Q_{i_0}\big(x_0,u_{i_0}(x_0),\partial_x\overline{\phi}_{i_0}(x_0),\partial_x^2\overline{\phi}_{i_0}(x_0)\big)\leq 0,~\text{if}~x_0 \in (0,R),\\
\min\Big\{\underset{i\in[I]}{\min}\{Q_{i}\big(0,u(0),\partial_x\overline{\phi}_{i}(0),\partial_x^2\overline{\phi}_{i}(0)\big)\},-F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial \overline{\phi}_I(0)\big)\Big\}\leq 0, \\
\text{if}~x_0 =\bf 0.
\end{cases}
\end{eqnarray*}
c) A generalized continuous viscosity Kirchhoff's solution is a continuous function on $\mathcal{N}_R$ that is simultaneously a generalized viscosity Kirchhoff's sub and super solution. 
\end{Definition}
\begin{Definition} \label{def : strong sur/sous solutions}\text{Viscous Kirchhoff's solutions: Ohavi \cite{Ohavi visco}}:\\
a) We say that $u\in \mathcal{L}_{sc}^b(\mathcal{N}_R,\R)$ is a viscosity Kirchhoff's super solution of
\eqref{eq PDE 0}, if for all test function $\overline{\phi} \in \mathcal{C}^{2}\big(\mathcal{N}_R,\R\big)$, and for all local minimum point $(x_0,i_0)\in \mathcal{N}_R$ of $u-\overline{\phi}$, with $u_{i_0}(x_0)-\overline{\phi}_{i_0}(x_0)=0$, we have:
\begin{eqnarray*}
\begin{cases}
Q_{i_0}\big(x_0,u_{i_0}(x_0),\partial_x\overline{\phi}_{i_0}(x_0),\partial_x^2\overline{\phi}_{i_0}(x_0)\big)\ge 0,~\text{if}~x_0 \in (0,R),\\
F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial_x \overline{\phi}_I(0)\big)\leq 0, ~~\text{if}~x_0 = \bf 0.
\end{cases}
\end{eqnarray*}
b) We say that $u\in \mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$ is a viscosity Kirchhoff's sub solution of
\eqref{eq PDE 0}, if for all test function $\overline{\phi} \in \mathcal{C}^{2}\big(\mathcal{N}_R,\R\big)$, and for all local maximum point $(x_0,i_0)\in \mathcal{N}_R$ of $u-\overline{\phi}$, with $u_{i_0}(x_0)-\overline{\phi}_{i_0}(x_0)=0$, we have:
\begin{eqnarray*}
\begin{cases}
Q_{i_0}\big(x_0,u_{i_0}(x_0),\partial_x\overline{\phi}_{i_0}(x_0),\partial_x^2\overline{\phi}_{i_0}(x_0)\big)\leq 0,~\text{if}~x_0 \in (0,R),\\
F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial_x \overline{\phi}_I(0)\big)\ge 0, ~~\text{if}~x_0 = \bf 0.
\end{cases}
\end{eqnarray*}
c) A viscosity continuous Kirchhoff's solution is a continuous function on $\mathcal{N}_R$ that is simultaneously a viscosity Kirchhoff's sub and super solution in the sense above. 
\end{Definition}
One of the main results of this work is the following Theorem that states that any viscosity generalized Kirchhoff's super/sub solution in the sense of Definition \ref{def : weak sur/sous solutions}, is indeed a viscosity Kirchhoff's super/sub solution given in Definition \ref{def : strong sur/sous solutions}.
\begin{Theorem}\label{th : weak are strong}
Assume assumption $(\mathcal{H})$. Let $u \in \mathcal{L}_{sc}^b(\mathcal{N}_R,\R)$ (resp. $v \in \mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$ be a viscosity generalized Kirchhoff's super (resp. sub) solution in the sense of Definition \ref{def : weak sur/sous solutions} of system \eqref{eq PDE 0}. Then $u$ (resp. $v$) is a viscosity Kirchhoff's super (resp. sub) solution, in the sense of Definition \ref{def : strong sur/sous solutions} of system \eqref{eq PDE 0}.   
\end{Theorem}
In the spirit of the recent advances obtained in \cite{Ohavi visco}, we will obtain the following comparison Theorem, without the  introduction of the deterministic "local-time" variable. Recall that we deal with the second order possibly degenerate framework, and then we are able to treat equations whose nature ranges from first-order Hamilton Jacobi Equations to uniformly elliptic second-order.
\begin{Theorem} (Comparison Theorem.)\label{th : compa syst}
Assume assumption $(\mathcal{H})$. Let $u \in \mathcal{L}_{sc}^b(\mathcal{N}_R,\R)$ be a Kirchhoff's super solution and  $v \in \mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$ be a Kirchhoff's sub solution of \eqref{eq PDE 0}. If the following boundary condition: 
$$\forall i\in [I],~~u_i(R)\ge v_i(R),$$
is satisfied, then:
$$\forall (x,i)\in \mathcal{N}_R,~~u_i(x)\ge v_i(x). $$
\end{Theorem}
Finally, combining the classical Perron's method and the last Theorem \ref{th : weak are strong}, we have:
\begin{Theorem} (Well-posedness) \label{th : existence}
Assume assumption $(\mathcal{H})$. Then there exists a unique continuous viscosity solution $u\in \mathcal{C}\big(\mathcal{N}_R,\R\big)$ of system \eqref{eq PDE 0}.
\end{Theorem}

\section{The First order case} \label{sec 1 ordre}
The aim of this Section is to prove the three main results Theorems \ref{th : weak are strong}, \ref{th : compa syst} and \ref{th : existence} in the case of first order problems for system \eqref{eq PDE 0}. More precisely, the system posed on the star-shaped network is the following one:
\begin{eqnarray}\label{eq : system first order}
\begin{cases}
Q_i\big(x,u_i(x),\partial_xu_i(x)\big)=0,~~x\in  (0,R),\\
F\big(u(0),\partial_xu_1(0),\ldots,\partial_xu_I(0)\big)=0,~~x=\bf 0,
\end{cases}.
\end{eqnarray}
Through this Section, we will work under assumption $(\mathcal{H})$. Note that c) is not required and b) is most naturally replaced by:
\begin{eqnarray}\label{eq assum rafinee}
\nonumber &\forall (x,y,u,p)\in[0,R]^2\times \R^2,~~\text{if}~~|u|\leq M,~~\exists C_M>0,\\
 &\text{s.t}~~\forall i\in[I]:~~|Q_i(x,u,p)-Q_i(y,u,p)|\leq C_M(1+|p|)|x-y|,  
\end{eqnarray}
in order to be able to use the doubling variable method inside each rays, and obtain a comparison theorem outside $\bf 0$.
We start with the following Proposition, which aims to build a precious test function for a super viscosity solution.
\begin{Proposition}\label{pr: const test sur sol}
Assume assumption $(\mathcal{H})$. Let $u\in \mathcal{L}_{sc}(\mathcal{N}_R,\R)$ and  $\mathcal{V}_\varepsilon=\bigcup_{i=1}^I[0,\varepsilon_i)$ a neighborhood of the junction point $\bf 0$ (where $\varepsilon=(\varepsilon_1,\ldots,\varepsilon_I)$, and $\forall i\in [I]$, $\varepsilon_i>0$ are small enough). Fix $(\theta,M,\lambda,C>0)$. Then there exist $\eta_i(\varepsilon)=\eta_i(\varepsilon,u,\lambda,C)>0,~i\in[I]$, such that the following ODE system posed on $\mathcal{V}_\varepsilon$:
\begin{eqnarray}\label{eq : system EDO test ordre 1}
\begin{cases}\lambda \overline{\phi}_i(x)-\lambda M+C\big(1+|\partial_x\overline{\phi}_i(x)|\big)=-\eta_i(\varepsilon),~~x\in  (0,\varepsilon_i),\\
\overline{\phi}_i(\varepsilon_i)-\overline{\phi}_i(0)=u_i(\varepsilon_i)-u(0)-\theta,\\
\forall (i,j)\in[I]^2,~~\overline{\phi}_i(0)=\overline{\phi}_j(0)=\overline{\phi}(0),
\end{cases},
\end{eqnarray}
admits a unique solution $\overline{\phi}=\overline{\phi}(\varepsilon,u,\lambda,C,\theta,M,(\eta_i(\varepsilon))_{i\in [i]}$ in the class $\mathcal{C}^1(\mathcal{V}_\varepsilon,\R)$.
\end{Proposition}
\begin{proof}
The idea is to proceed by induction and to solve edge by edge each ordinary differential equation, adjusting  at each step the parameters $(\eta_i(\varepsilon))_{i\in [i]}$, in order to guarantee the continuity of the solution $\overline{\phi}$ at the junction point, that reads:
$$\forall (i,j)\in[I]^2,~~\overline{\phi}_i(0)=\overline{\phi}_j(0)=\overline{\phi}(0).$$
At each step $i\in [I]$, depending on the sign of $u_i(\varepsilon_i)-u(0)-\theta$, it is easy to check that the solution of:
\begin{eqnarray}\label{edo 1}
\nonumber \lambda\overline{\phi}_i(x)-\lambda M+C\big(1+|\partial_x\overline{\phi}_i(x)|\big)=-\eta_i(\varepsilon),~~x\in  (0,\varepsilon_i),\\
\overline{\phi}_i(\varepsilon)-\overline{\phi}_i(0)=u_1(\varepsilon_i)-u(0)-\theta,  
\end{eqnarray}
is given $\forall x\in [0,\varepsilon_i]$ by:
\begin{eqnarray}\label{expression direct function test}
\nonumber&\overline{\phi}_i(x)=\\
&\begin{cases}\ds \frac{u_i(\varepsilon_i)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_i)-1}\exp(-\frac{\lambda}{C}x)-\frac{C+\eta_i(\varepsilon)}{\lambda}+M,~~&\text{if}~u_i(\varepsilon_i)-u(0)-\theta \ge 0,\\
\ds \frac{u(\varepsilon_i)-u(0)-\theta}{\exp(\frac{\lambda}{C}\varepsilon_i)-1}\exp(\frac{\lambda}{C}x)-\frac{C+\eta_i(\varepsilon)}{\lambda}+M,~~&\text{if}~u_i(\varepsilon_i)-u(0)-\theta\leq 0.
\end{cases}
\end{eqnarray}
Hence, if for example $u_1(\varepsilon_1)-u(0)-\theta \ge 0$ and $u_2(\varepsilon_2)-u(0)-\theta \ge 0$, to obtain at the first step that $\phi_1(0)=\phi_2(0)$,  $\eta_1(\varepsilon)>0$ and $\eta_2(\varepsilon)>0$ should satisfy:
\begin{eqnarray}\label{eq equality}
\frac{u_1(\varepsilon_1)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_1)-1}-\frac{C+\eta_1(\varepsilon)}{\lambda}+M=\frac{u_2(\varepsilon_2)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_2)-1}-\frac{C+\eta_2(\varepsilon)}{\lambda}+M.\end{eqnarray}
If $\eta_1(\varepsilon)>0$ is then fixed first and verifies:
$$\eta_1(\varepsilon)>\max\Big(\lambda \Big[\frac{u_2(\varepsilon_2)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_2)-1}-\frac{u_1(\varepsilon_1)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_1)-1}\Big],0\Big),$$
we can choose $\eta_2(\varepsilon)>0$ to obtain equality \eqref{eq equality}. The other cases involving different signs of $u_1(\varepsilon_1)-u(0)-\theta$ and $u_2(\varepsilon_2)-u(0)-\theta$ can be treated similarly to obtain that $\overline{\phi}_1(0)=\overline{\phi}_2(0)$ as soon as $\eta_1(\varepsilon)$ is large enough. 
Finally, it is easy to see that we will obtain the existence of the sequence $(\eta_i(\varepsilon))_{i \in [I]}$ that will guarantee the continuity of the solution $\overline{\phi}$ of \eqref{eq : system EDO test ordre 1} at the junction point $\bf 0$: 
$$\forall (i,j)\in[I]^2,~~\overline{\phi}_i(0)=\overline{\phi}_j(0)=\overline{\phi}(0),$$
repeating the last arguments step by step.
\end{proof}
\begin{Proposition}\label{pr: fonction test est forte}
Assume assumption $(\mathcal{H})$. Let $u\in \mathcal{L}^b_{sc}(\mathcal{N}_R,\R)$ a generalized Kirchhoff super solution of system \eqref{eq : system first order}. Then the associated map $\overline{\phi}\in \mathcal{C}^1(\mathcal{V}_\varepsilon,\R)$ constructed in Proposition \ref{pr: const test sur sol}, is a test function of $u$ and at the junction point $\bf 0$ and satisfies the Kirchhoff's boundary condition:
\begin{eqnarray}\label{eq Kirchoff uper phi}
F\big(u(0),\partial_x\overline{\phi}_1(0),\ldots,\partial_x{\phi}_I(0)\big)\leq 0.   
\end{eqnarray}
\end{Proposition}
\begin{proof}
By assumption, there exists $M_u>0$ such that:
$$\forall i \in [I],~~\forall x\in [0,R],~~u_i(x)<M_u.$$
Let $\overline{\phi}=\overline{\phi}(\varepsilon,u,M_u,\lambda,C_{M_u},\theta,(\eta_i(\varepsilon))_{i\in [i]})\in \mathcal{C}^1(\mathcal{V}_\varepsilon,\R)$ be the associated function of $u$ constructed in Proposition \ref{pr: const test sur sol}, satisfying \eqref{eq : system EDO test ordre 1} (where the constants $(\lambda,C_{M_u})$ appear in assumption $(\mathcal{H}$ a)-d)).
To obtain the result, we need to show that the minimum of $u-\overline{\phi}$ can be reached only at the junction point $\bf 0$. Indeed, remark first by definition of $\overline{\phi}$ in \eqref{eq : system EDO test ordre 1}, that we have:
$$\forall i\in [I],~~u(0)-\phi(0)=u_i(\varepsilon_i)-\phi_i(\varepsilon_i)-\theta < u_i(\varepsilon_i)-\phi_i(\varepsilon_i), $$
(recall $\theta>0$), implying that the minimum can not be reached at the boundaries $(\varepsilon_i)_{i \in [I]}$. On the other hand, if the minimum of $u-\overline{\phi}$ is reached for some $y\in (0,\varepsilon_j)$, for a given $j\in [I]$, since $u$ is a super solution (after a standard modification to obtain $\overline{\phi}_j(y)=u_j(y)$), we have:
\begin{eqnarray}\label{eq sur sol 1}
Q_j\big(y,u_j(y),\partial_x\overline{\phi}_j(y)\big)\ge 0.
\end{eqnarray}
On the other hand, from assumptions $(\mathcal{H})-d)$ and \eqref{eq : system EDO test ordre 1}, we obtain:
\begin{align}\label{eq maj eikonal}
\nonumber Q_j\big(y,u_j(y),\partial_x\overline{\phi}_j(y)\big) \leq \lambda (u_j(y)-M_u)+Q_j\big(y,M_u,\partial_x\overline{\phi}_j(y)\big)\\
\leq \lambda \overline{\phi}_j(y)-\lambda M_u+C_{M_u}(1+|\partial_x\overline{\phi}_j(y)|)=-\eta_j(\varepsilon)<0, \end{align}
and then contradicts \eqref{eq sur sol 1}. We conclude therefore that the minimum of $u-\overline{\phi}$ is reached at the junction point $\bf 0$, and since $u$ is a generalized Kirchhoff's super solution, we obtain:
$$\max\Big\{~\underset{i\in[I]}{\max}\big\{~Q_{i}\big(0,u(0),\partial_x\overline{\phi}_{i}(0)\big)~\big\},-F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial \overline{\phi}_I(0)\big)~\Big\}\ge 0.$$
But since $\forall i\in [I]$ (once again with the aid of \eqref{eq : system EDO test ordre 1}):
\begin{align*}
Q_{i}\big(0,u(0),\partial_x\overline{\phi}_{i}(0)\big)) \leq \lambda (u(0)-M_u)+Q_i\big(0,M_u,\partial_x\overline{\phi}_i(0)\big)\\
\leq \lambda \phi(0)-\lambda M_u+C_{M_u}(1+|\partial_x\overline{\phi}_i(0)|)=-\eta_i(\varepsilon)<0, \end{align*}
we have that necessary:
$$F\big(u(0),\partial_x \overline{\phi}_1(0),\ldots,\partial_x \overline{\phi}_I(0)\big)\leq 0,$$
and that completes the proof.
\end{proof}
One of the central results of this work, is to show that any generalized Kirchhoff's super/sub viscosity solution (in the weak sense Definition \ref{def : weak sur/sous solutions}) is indeed a Kirchhoff's super/sub viscosity solution (in the strong sense, Definition \ref{def : strong sur/sous solutions}). That is the concern of the next Proposition for the first order (in the super solution case).
\begin{Proposition}\label{pr: weak is strong}
Assume assumption $(\mathcal{H})$. Let $u\in \mathcal{L}^b_{sc}(\mathcal{N}_R,\R)$ be a generalized viscous Kirchhoff's super solution of system \eqref{eq : system first order}. Then $u$ is a viscous Kirchhoff's super solution of system \eqref{eq : system first order}. 
\end{Proposition}
\begin{proof}
To obtain such a statement, we consider $\psi$ a test function of $u$ at the junction point $\bf 0$ and we will show that: 
$$F\big(u(0),\partial_x \psi_1(0),\ldots,\partial_x \psi_I(0)\big)\leq 0.$$
There exists a neighborhood of the junction $\mathcal{V}_\kappa$ that we will denote for the sake of simplicity:
$$\mathcal{V}_\kappa=\bigcup_{i=1}^I[0,\kappa),$$
with $\kappa>0$ small enough, such that:
$$\underset{(x,i)\in \overline{\mathcal{V}_\kappa}}{\min}\{u_i(x)-\psi_i(x)\}=u(0)-\psi(0)=0.$$
For any $\varepsilon>0$ small enough, ($\varepsilon<<\kappa$), we introduce: $$\overline{\phi}=\overline{\phi}^{\varepsilon,\theta}=\overline{\phi}(\varepsilon,u,M_u,\lambda,C_{M_u},\theta,(\eta_i(\varepsilon))_{i\in [i]})\in \mathcal{C}^1(\mathcal{V}_\varepsilon,\R),$$ 
the associated test function of $u$ at $\bf 0$, constructed in Proposition \ref{pr: const test sur sol}, satisfying \eqref{eq : system EDO test ordre 1} (note that $\varepsilon_i=\varepsilon,~\forall i\in [I]$, $\theta>0$, $u<M_u$ and the constants $(\lambda,C_{M_u})$ appear in assumption $(\mathcal{H})$). Because $\psi$ is a test function of $u$ at $\bf 0$, first remark that for all $i\in [I]$:
$$u(0)-\psi(0)\leq u_i(\varepsilon)-\psi_i(\varepsilon),~~\text{then}~~\psi_i(\varepsilon)-\psi(0)\leq u_i(\varepsilon)-u(0).$$
Hence using the expression of the boundary condition of $\overline{\phi}^{\varepsilon,\theta}$ given in \eqref{eq : system EDO test ordre 1}, we obtain that for all $i\in [I]$:
$$\psi_i(\varepsilon)-\psi(0)\leq u_i(\varepsilon)-u(0)=\overline{\phi}^{\varepsilon,\theta}_i(\varepsilon)-\overline{\phi}^{\varepsilon,\theta}(0)+\theta.$$
Now, using the expression of $\overline{\phi}^{\varepsilon,\theta}$ given in \eqref{expression direct function test}, we get that $x\mapsto \partial_x\phi_i(x)$ is decreasing $\forall i \in [I]$, leading therefore to:
\begin{eqnarray}\label{eq key 0}
\psi_i(\varepsilon)-\psi(0)\leq \int_0^\varepsilon\partial_x\overline{\phi}_i^{\varepsilon,\theta}(u)du+\theta\leq \varepsilon\partial_x \overline{\phi}^{\varepsilon,\theta}_i(0)+\theta,    
\end{eqnarray}
and then: $$\frac{\psi_i(\varepsilon)-\psi(0)}{\varepsilon}-\frac{\theta}{\varepsilon}\leq \partial_x \overline{\phi}^{\varepsilon,\theta}_i(0),~~\forall i\in[I]. $$
From assumption $(\mathcal{H})$, $F$ is increasing with respect to the gradient variable, and this implies using \eqref{eq Kirchoff uper phi}:
\begin{eqnarray*}
&\ds F\big(u(0),\frac{\psi_1(\varepsilon)-\psi(0)}{\varepsilon}-\frac{\theta}{\varepsilon},\ldots,\frac{\psi_I(\varepsilon)-\psi(0)}{\varepsilon}-\frac{\theta}{\varepsilon})\leq\\
&F\big(u(0),\partial_x \overline{\phi}^{\varepsilon,\theta}_1(0),\ldots,\partial_x \overline{\phi}^{\varepsilon,\theta}_I(0)) \leq 0.    
\end{eqnarray*}
Sending first $\theta\searrow 0$ and thereafter $\varepsilon\searrow 0$ in the last inequality, using the continuity of $F$, it follows that: 
\begin{eqnarray*}
&F\big(u(0),\partial_x \psi_1(0),\ldots,\partial_x \psi_I(0))=\\
&\ds \lim_{\varepsilon \searrow 0}\overline{\lim}_{\theta\searrow 0}F\big(u(0),\frac{\psi_1(\varepsilon)-\psi(0)}{\varepsilon}-\frac{\theta}{\varepsilon},\ldots,\frac{\psi_I(\varepsilon)-\psi(0)}{\varepsilon}-\frac{\theta}{\varepsilon})\leq 0.    
\end{eqnarray*}
and that completes the proof.
\end{proof}
\begin{Remark}\label{rm: constru test sub sol}
For the sub solution case, assuming assumption $(\mathcal{H})$, and given $v\in \mathcal{U}_{sc}^b(\mathcal{N}_R,\R)$ a generalized viscous Kirchhoff's sub solution of system \eqref{eq : system first order}, such that $v >-M_v$, repeating the same arguments of Proposition \ref{pr: const test sur sol}, we will introduce naturally the associated map $\underline{\phi}=\underline{\phi}(\varepsilon,u,M_v,\lambda,C_{M_v},\theta,(\eta_i(\varepsilon))_{i\in [i]})\in \mathcal{C}^1(\mathcal{V}_\varepsilon,\R)$, solution of the following system:
\begin{eqnarray}\label{eq : system EDO test sous ordre 1}
\begin{cases}\lambda \underline{\phi}_i(x)+\lambda M_v-C_{M_v}\big(1+|\partial_x\underline{\phi}_i(x)|\big)=\eta_i(\varepsilon),~~x\in  (0,\varepsilon_i),\\
\underline{\phi}_i(\varepsilon_i)-\underline{\phi}_i(0)=v_i(\varepsilon_i)-v(0)+\theta,\\
\forall (i,j)\in[I]^2,~~\underline{\phi}_i(0)=\underline{\phi}_j(0)=\underline{\phi}(0),
\end{cases},
\end{eqnarray}
by guaranteeing that the constants $(\eta_i(\varepsilon))_{i\in [i]}$ are chosen strictly positive. The explicit expression of $\underline{\phi}$ is given by: 
\begin{eqnarray}\label{expression direct function test sub}
\nonumber&\underline{\phi}_i(x)=\\
&\begin{cases}\ds \frac{v_i(\varepsilon_i)-v(0)+\theta}{\exp(\frac{\lambda}{C_{M_v}}\varepsilon_i)-1}\exp(\frac{\lambda}{C_{M_v}}x)+\frac{C_{M_v}+\eta_i(\varepsilon)}{\lambda}-M_v,~~&\text{if}~v_i(\varepsilon_i)-v(0)+\theta \ge 0,\\
\ds \frac{v_i(\varepsilon_i)-v(0)+\theta}{\exp(-\frac{\lambda}{C_{M_v}}\varepsilon_i)-1}\exp(-\frac{\lambda}{C_{M_v}}x)+\frac{C_{M_v}+\eta_i(\varepsilon)}{\lambda}-M_v,~~&\text{if}~v_i(\varepsilon_i)-v(0)+\theta\leq 0.
\end{cases}
\end{eqnarray}
Repeating the same arguments given in Proposition \ref{pr: fonction test est forte}, it is easy to check that $\underline{\phi}$ is a test function at $\bf 0$, for the sub solution $v$, and satisfies:
\begin{eqnarray}\label{eq Kirchoff sub phi}
F\big(v(0),\partial_x\underline{\phi}_1(0),\ldots,\partial_x\underline{\phi}_I(0)\big)\ge 0.   
\end{eqnarray}
Once again, the same arguments given in Proposition \ref{pr: weak is strong}, will imply that any $v\in \mathcal{U}^b_{sc}(\mathcal{N}_R,\R)$  generalized viscous Kirchhoff's sub solution of system \eqref{eq : system first order}, will be in fact a viscous Kirchhoff's sub solution of system \eqref{eq : system first order}. 
The last results and discussions then allow us to state that Theorem \ref{th : weak are strong} holds true for first order systems of type \eqref{eq : system first order}.
\end{Remark}
Next we prove a comparison theorem.

\textbf{Proof of the comparison Theorem \ref{th : compa syst} for system of first order of type \eqref{eq : system first order}}.
\begin{proof}
As it is classical in viscosity solution posed on networks, the main difficulties arise at the junction point $\bf 0$. Indeed, if we assume that the following supremum $\sup_{(x,i)\in \mathcal{N}_R}=\{v_i(x)-u_i(x)\}>0$ is reached outside the junction point, we will obtain a contradiction using classical arguments of doubling variable method and assumption $(\mathcal{H})$. Hence, we will only show that the following assumption
\begin{eqnarray}\label{assumpt contra}
    \sup_{(x,i)\in \mathcal{N}_R} \{v_i(x)-u_i(x)\}=v(0)-u(0)>0,
\end{eqnarray}
will lead to a contradiction. 
Using the two test functions built at the junction point $\bf 0$ for $u$ and $v$ in Proposition \ref{pr: fonction test est forte} and Remark \ref{rm: constru test sub sol} (with the same strict upper/lower bound $M>0$ for $u$ and $v$), it follows that we have:
$$F\big(u(0),\partial_x \overline{\phi}^{\varepsilon,\theta}_1(0),\ldots,\partial_x \overline{\phi}^{\varepsilon,\theta}_I(0))\leq 0,~~F\big(v(0),\partial_x \underline{\phi}^{\varepsilon,\theta}_1(0),\ldots,\partial_x \underline{\phi}^{\varepsilon,\theta}_I(0))\ge 0,$$
It is easy to check with both expressions of $\overline{\phi}^{\varepsilon,\theta}$ and $\underline{\phi}^{\varepsilon,\theta}$ given in \ref{expression direct function test} and \eqref{expression direct function test sub}, that for any $\varepsilon>0$, the last two functions will converge uniformly in $\mathcal{C}^1(\mathcal{V}_\varepsilon)$ as $\theta \searrow$, to $\overline{\phi}^{\varepsilon}$ and $\underline{\phi}^{\varepsilon}$ such that 
$$F\big(u(0),\partial_x \overline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \overline{\phi}^{\varepsilon}_I(0))\leq 0,~~F\big(v(0),\partial_x \underline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \underline{\phi}^{\varepsilon}_I(0))\ge 0.$$
Therefore, we get:
\begin{eqnarray}\label{eq F sub super}
F\big(u(0),\partial_x \overline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \overline{\phi}^{\varepsilon}_I(0))\leq F\big(v(0),\partial_x \underline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \underline{\phi}^{\varepsilon}_I(0)).\end{eqnarray}
Now remark that there exists $j\in [I]$ such that:
\begin{eqnarray}\label{eq contra grad}
\partial_x \overline{\phi}^{\varepsilon}_j(0)\leq \partial_x \underline{\phi}^{\varepsilon}_j(0),  
\end{eqnarray}
since, if one assumes the contrary, assumption $(\mathcal{H}-e)-f))$ implies that $F$ is strictly increasing with respect to its gradient and decreasing with respect to its first variable, and then (recall that $v(0)\ge u(0)$): 
$$F\big(u(0),\partial_x \overline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \overline{\phi}^{\varepsilon}_I(0))>F\big(v(0),\partial_x \underline{\phi}^{\varepsilon}_1(0),\ldots,\partial_x \underline{\phi}^{\varepsilon}_I(0)),$$
that is a contradiction with \eqref{eq F sub super}.
Now we claim that:
$$u_j(\varepsilon_j)-u(0)=v_j(\varepsilon_j)-v(0).$$
Recall that \eqref{assumpt contra} implies $u_j(\varepsilon_j)-u(0)\ge v_j(\varepsilon_j)-v(0).$ Moreover, we have:
$$\partial_x\overline{\phi}^\varepsilon_j(0)=\\
\begin{cases}\ds \frac{\lambda}{C}\frac{u_j(\varepsilon_j)-u(0)}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)}&\text{if}~u_i(\varepsilon_j)-u(0)\ge 0,\\
\ds \frac{\lambda}{C}\frac{u(\varepsilon_i)-u(0)-\theta}{\exp(\frac{\lambda}{C}\varepsilon_j)-1},~~&\text{if}~u_j(\varepsilon_j)-u(0)\leq 0,
\end{cases}
$$
and:
$$
\partial_x\underline{\phi}_j^\varepsilon(0)=\\
\begin{cases}\ds \frac{\lambda}{C}\frac{v_j(\varepsilon_j)-v(0)}{\exp(\frac{\lambda}{C}\varepsilon_j)-1},~~&\text{if}~v_j(\varepsilon_j)-v(0) \ge 0,\\
\ds \frac{\lambda}{C}\frac{v_j(\varepsilon_j)-v(0)+\theta}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)},~~&\text{if}~v_j(\varepsilon_j)-v(0)\leq 0.
\end{cases}
$$
We argue in the next lines by contradiction assuming that: $u_j(\varepsilon_j)-u(0)> v_j(\varepsilon_j)-v(0).$ It is also easy to be convinced that 
$$\frac{1}{\exp(b)-1}\leq\frac{1}{1-\exp(-b)},~~\forall b>0.$$
We distinguish then only three cases: if $u_j(\varepsilon_j)-u(0)>v_j(\varepsilon_j)-v(0)>0$, then:
$$\partial_x\overline{\phi}_j(0)=\frac{\lambda}{C}\frac{u_j(\varepsilon_j)-u(0)}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)}>\frac{\lambda}{C}\frac{v_j(\varepsilon_j)-v(0)}{\exp(\frac{\lambda}{C}\varepsilon_j)-1}=\partial_x\underline{\phi}_j(0),$$
if $u_j(\varepsilon_j)-u(0)>0>v_j(\varepsilon_j)-v(0)>0$, then:
$$\partial_x\overline{\phi}_j(0)=\frac{\lambda}{C}\frac{u_j(\varepsilon_j)-u(0)}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)}>\frac{\lambda}{C}\frac{v_j(\varepsilon_j)-v(0)}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)}=\partial_x\underline{\phi}_j(0),$$
if $0\ge u_j(\varepsilon_j)-u(0)>v_j(\varepsilon_j)-v(0)>0$, then:
$$\partial_x\overline{\phi}_j(0)=\frac{\lambda}{C}\frac{u_j(\varepsilon_j)-u(0)}{\exp(\frac{\lambda}{C}\varepsilon_j)-1}>\frac{\lambda}{C}\frac{v_j(\varepsilon_j)-v(0)}{1-\exp(-\frac{\lambda}{C}\varepsilon_j)}=\partial_x\underline{\phi}_j(0).$$
In other words, all the three last cases contradict \eqref{eq contra grad}. We can conclude therefore that:
$$u_j(\varepsilon_j)-u(0)=v_j(\varepsilon_j)-v(0).$$
In conclusion, the supremum $\sup_{(x,i)\in \mathcal{N}_R}=\{v_i(x)-u_i(x)\}$ is reached outside the junction point, and this will lead to a contradiction. The proof is complete.
\end{proof}
Now we are able to prove Theorem \ref{th : existence} in the case of first order systems of type \eqref{eq : system first order} under assumption $(\mathcal{H})$.
\begin{proof}
\textbf{Uniqueness:} It is a direct consequence of the last comparison Theorem \ref{th : compa syst} for system \eqref{eq : system first order}. \\
\textbf{Existence:} We start by showing that the two following maps defined on $\mathcal{N}_R$, by: $(x,i)\mapsto A+B\exp(-x)$ and $(x,i)\mapsto-A-B\exp(-x)$ are super and sub Kirchhoff's viscous solutions of system \eqref{eq : system first order}, as soon as we have:
$$B\ge\frac{|F(0,0,\ldots,0)|}{I\alpha},~\text{and}~A\ge \frac{1}{\lambda}\max_i|Q_i(x,0,-B\exp(-x))|_{[0,R]}.$$
For the super solution case, let us check the calculations. We have from assumption $(\mathcal{H}) -e) -f)$ at $\bf 0$:
$$F(A+B,-B,\ldots-B)\leq F(0,-B,\ldots-B)\leq F(0,0,\ldots,0)-I\alpha B \leq 0.$$
Assumption $(\mathcal{H}) -a)$, with the choice of $A$ implies $\forall i\in [I],~x\in (0,R)$:
\begin{eqnarray*}
 &Q_i(x,A+B\exp(-x),-B)\ge \lambda A + Q_i(x,B\exp(-x),-B\exp(-x))\\
 &\ge \lambda A + Q_i(x,0,-B\exp(-x))\ge 0.   
\end{eqnarray*}
Same arguments hold true to show that $(x,i)\mapsto-A-B\exp(-x)$ is sub Kirchhoff's viscous of solutions of \eqref{eq : system first order}.
We can then use the classical Perron's method and the last comparison Theorem \ref{th : compa syst}, to obtain the existence of a generalized Kirchhoff's viscous continuous viscosity solution for system \eqref{eq : system first order} (in the sense of Definition \ref{def : weak sur/sous solutions}). We conclude using that any generalized Kirchhoff's continuous viscous solution is a Kirchhoff's continuous viscous solution (Theorem \ref{th : weak are strong}).
\end{proof}
\section{The Second Order Case}\label{sec 2 order}
In this Section, we will see that Theorems \ref{th : weak are strong}-\ref{th : compa syst}-\ref{th : existence} hold true, under assumption $(\mathcal{H})$, for the second order system \eqref{eq PDE 0}, in the possibly degenerate case. Indeed, we will see that we can easily adapt the techniques of proof that we have introduced in the last Section \ref{sec 1 ordre} for the first order. Since the second and first order systems \eqref{eq PDE 0}-\eqref{eq : system first order}, have exactly the same boundary condition $F$ at $\bf 0$, it is then natural to ask if both test functions $(\overline{\phi},\underline{\phi})$ built in Proposition \ref{pr: const test sur sol} and Remark \ref{rm: constru test sub sol}, are also good test functions candidates for the second order case involving \eqref{eq PDE 0}. For that purpose, in order to be able to obtain an inequality of the type \eqref{eq maj eikonal}, in the super solution case with the aid of assumption $(\mathcal{H}-d)$, we have to control uniformly the bound of the second order derivatives of $\overline{\phi}$. Recall now that from \eqref{expression direct function test}, we have:
\begin{eqnarray*}
\nonumber&\forall i\in [I],~~\forall x\in [0,\varepsilon_i],~~\partial_x^2\phi_i(x)=\\
&\begin{cases}\ds \big(\frac{\lambda}{C}\big)^2\frac{u_i(\varepsilon_i)-u(0)-\theta}{\exp(-\frac{\lambda}{C}\varepsilon_i)-1}\exp(-\frac{\lambda}{C}x),~~&\text{if}~u_i(\varepsilon_i)-u(0)-\theta \ge 0,\\
\ds \big(\frac{\lambda}{C}\big)^2\frac{u(\varepsilon_i)-u(0)-\theta}{\exp(\frac{\lambda}{C}\varepsilon_i)-1}\exp(\frac{\lambda}{C}x),~~&\text{if}~u_i(\varepsilon_i)-u(0)-\theta\leq 0.
\end{cases}
\end{eqnarray*}
Hence if $M>|u|$, we get that $\forall i\in [I],~~\forall x\in [0,\varepsilon_i]$:
$$-\frac{2(M+\theta)\lambda}{C\varepsilon}\exp(\frac{\lambda\varepsilon}{C})\leq \partial_x^2\overline{\phi}_i(x)\leq 0.$$
Since the constant $C>0$ can be chosen large enough in $(\mathcal{H}-d)$ to obtain \eqref{eq maj eikonal}, for example $C_\varepsilon>>\frac{1}{\varepsilon}\vee C$, we see that we can impose in that case an uniform lower bound for $\partial_x^2\overline{\phi}$, and for instance: $-1\leq \partial_x^2\overline{\phi}\leq 0$, (recall $\theta <<1$). It follows then, in the spirit of \eqref{eq maj eikonal} that $\forall i\in [I],~~\forall x\in [0,\varepsilon_i]$:
\begin{align*}
&Q_i\big(x,u_i(x),\partial_x\overline{\phi}_i(x),\partial_x^2\overline{\phi}_j(x)\big)
\leq \lambda (u_i(x)-M)+Q_i\big(x,M,\partial_x\overline{\phi}_i(x),-1\big)\\
&\leq \lambda u_i(x)-\lambda M+C_{M,-1}(1+|\partial_x\overline{\phi}_i(x)|)=\lambda \phi_i(x)-\lambda M+C_{\varepsilon}(1+|\partial_x\overline{\phi}_i(x)|),\end{align*}
(where we have denoted $C_{M,-1}=C$). We treat similarly the sub solution case to build a test function at $\bf 0$ for the second order system \eqref{eq PDE 0}, controlling uniformly the upper bound of its second derivative. It follows that all the arguments and techniques of proof of the last Section \ref{sec 1 ordre} will hold true, in the second order possibly degenerated case. The last discussion allows us then to conclude the validity of Theorems \ref{th : weak are strong}-\ref{th : compa syst}-\ref{th : existence}.

\end{document}